OpenStudy (sarahc):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -3 + 3 sin θ y-axis only Origin only x-axis only No symmetry

OpenStudy (jhannybean):

Think of the coordinate points \((r~,~\theta)\). \[(r,\theta) \rightarrow (-r,~\pi -\theta) \implies \text{symmetric about x-axis}\]\[(r,\theta) \rightarrow (r,~\pi -\theta) \implies \text{symmetric about the y-axis} \]\[(r,\theta) \rightarrow (r,~ \theta +\pi ) \implies \text{symmetric about the origin}\] plugging in or analyzing coordinate points for these functions will usually tell you whether or not your function is symmetric, and about which axis or origin.

OpenStudy (rational):